Factorise the expressions and divide them as directed. 
(y^2 +7y+10)÷(y+5)

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

First solve for equation, (y^2+7y+10)
(y^2+7y+10) = y^2+2y+5y+10
= y(y+2)+5(y+2) = (y+2)(y+5)
Now, (y^2+7y+10)÷(y+5)
= (y+2)(y+5)/(y+5) = y+2

Thanks